# How Consistent Are Humans When Grading Programming Assignments?

### v1

@@ 4.5. Grading Consistency of Multiple Graders (RQ1)
Krippendorff’s $\alpha$ is a percent chance-based reliability coefficient, similar to Cohen’s $\kappa$ and Fleiss’ $\kappa$, and evaluates the deviation from perfect reliability by the proportion of observed to expected disagreement (Krippendorff, 2013).
@@
Figure 16 shows the $\alpha$ for each of the seven groups and each skill. These results show that in our study, consistency was low, but were slightly more likely to agree on the grades for correctness and code elegance but did not agree on the grades for readability or documentation, with the average $\alpha$ being 0.220210, 0.098628, 0.045598, and 0.054779, respectively.
Many percent chance agreement-based coefficients, including Krippendorf’s $\alpha$ and Fleiss’ $\kappa$, underestimate the agreement when one of the labels is dominant (Gwet, 2008; Doughty et al., 2024; Wongpakaran et al., 2013). As such, we compute Gwet’s $AC_{2}$, which is a weighted variant to Gwet’s $AC_{1}$, which was designed to be robust in the high prevalence of a single class (Gwet, 2008). Gwet’s $AC_{2}$ is more robust in instances of high prevalence, as it removes from consideration all agreements that occurred by chance (Gwet, 2008).
Figure 16 shows the results of Gwet’s $AC_{2}$. Following the same bounds defined by Krippendorf, no groups achieve acceptable agreement ($AC_{2}>0.800$) for any of the skills evaluated. However, some groups show partial agreement with $AC_{2}\geq 0.667$ for specific skills, though on average, no skill had an $AC_{2}\geq 0.667$, with the average $AC_{2}$ for correctness being 0.628587, for code elegance being 0.592377, for readability being 0.638274 and for documentation being 0.596593.
To further explore the inter-rater reliability, we evaluated the spread of grades awarded by each group. Figure 17 shows the minimum, maximum, and mean correctness grades for group 17-20, which was the most consistent when grading correctness, with an $\alpha=0.45$, and for group 5-8, which was the least consistent when grading correctness, with an $\alpha=-0.069$. Appendix B shows the distribution of grades for all groups and all grades.
@@
Figure 17. These figures are for the highest and lowest $\alpha$ and show the minimum, maximum and mean correctness grade awarded by the participants for each assignment. Plots for all groups and all grades can be found in Appendix B.
To further investigate the reliability when awarding each grade, we stipulate that assignments grade ranges less than or equal to two (e.g. A to B+ is two grades away, and A to A- is one grade away) have good reliability; those with a range of three (e.g. A+ to B+) have mediocre reliability, and those with a range greater than three (e.g. A- to C+) have poor reliability; where the grade range is the maximum awarded grade minus the minimum awarded grade for each assignment. Figure 18 shows the proportion of the reliability for each grade and skill.
@@
4.6. Grading Consistency of Individual Graders (RQ2)
Only two participants, three and ten, provided the same grade in both batches for all four skills. Participant 10 noticed the duplicate, and during the interview, participant 10 mentioned that they noticed the duplicate and changed the grades for both batches to F as they believed it was plagiarism, while the other five participants followed the instruction that if you notice a duplicate, treat it as you are grading it the first time; as this was a pair programming assignment, duplicates could still exist within the dataset. Participant 3 did not notice the duplicate and said, “…like some assignments used the same animals… but [the implementation] was a little bit different so not like an exact copy of those two assignments…”.
Figure 21 shows the distribution of the absolute distance between grades awarded in batch one and batch two for each of the four skills. Those with no prior grading experience have the largest difference between grades across all skills. Those with 1-2 years of grading experience have at most a difference of three grades between the first and second time they graded the duplicated assignment, and those with one or more years of grading have a grade difference of less than one for most skills. However, readability and documentation for those with two or three years of grading experience had a maximum grade distance of two or three, respectively.
5. Discussion
5.1. How Consistently Does a Group of Graders Apply a Rubric? (RQ1)
Consistency was very poor overall. Even the best values for Krippendorf’s $\alpha$ achieved reflect distinctly poor consistency, with none of our seven groups of graders achieving an $\alpha>0.667$, the lower bound to consider reliable (Krippendorff, 2013), with the average $\alpha=0.10480$ across all groups and skills. Out of the four skills that the participants graded, correctness was most likely to be graded more consistently, with five of the seven groups achieving a higher $\alpha$ than the other four skills, with groups 5-8 and 25-28 achieving the second most consistently graded skill for their respective groups. In comparison, all participant groups were similarly inconsistent when grading code elegance, readability and documentation, all skills achieving an average $\alpha<0.1$.
The correctness of a program is, in most cases, inherently objective, though there can be elements of subjectiveness, especially within open-ended coursework assignments. The nature of open-ended assignments, where students decide how they implement a set of features using their ideas and creativity, can make it difficult for the groups of assessors to grade consistently.
Our most consistent group while grading correctness was 17-20, with an $\alpha=0.45$; Figure 17(a) shows the range of the awarded grades for the 40 assignments that they graded. Only assignments awarded an F (assignments 80 and 98) – since they either did not compile or did not include any student code – were assessed with complete agreement. The original authors of assignment 80 submitted the template code they provided, and assignment 98 did not compile or, if the assessors fixed the compilation error, had an unhandled NumberFormatException, causing the simulator not to execute.
Group 17-20, which had the highest average $\alpha=0.2571$ of three master’s students and one third-year student. Another group with a similar composition was group 1-4, consisting of three master’s students and one PhD student; however, they had an average $\alpha=0.1225$, the fourth lowest average $\alpha$.
Our participants follow the typical demographics of teaching assistants, being undergraduate, master’s or PhD students (Mirza et al., 2019), and many are new to assessment and have little experience. None of our participants met the threshold for partial agreement ($\alpha>0.667$). While having experienced teaching assistants to conduct the grading could improve the consistency, having those with many years of grading experience is typically rare; at our institution, many teaching assistants do not start assessing student work until they have nearly finished their degree. However, those with one year of grading experience had similar inter-rater reliability to those with no previous grading experience, indicating that more grading experience alone will not improve the grading consistency to an acceptable level.
@@ 5.1.1. An In-Depth Analysis of the Grades Awarded by Group 17-20
This trend of a few assignments being somewhat consistently graded and many being inconsistently graded holds for readability and documentation, with Appendix B showing the grade range for each skill. Figure 18 shows this trend for each grade, with the middle grades (D - B+) more likely to have a larger variance in awarded grades across all four skills. While the middle grades showed no cases of any grading achieving acceptable levels of consistency, the high (A++ - A) and low ranges (F - D-) displayed at least some instances of grader agreement. Grades F and D- have the largest portions of grade ranges, being less than or equal to two. At the other end of the grade spectrum, grades A, A+ and A++ have a smaller proportion of grade ranges that are less than or equal to two.
Submissions at the failing end of the grade spectrum may make it easier to provide consistent grading, as they are more likely to be incorrect, uncompilable, poorly designed, incomprehensible or missing documentation. Meanwhile, grades at the opposite end of the grade spectrum may make it easier to provide consistent grading. They are more likely to have fully-featured code that compiles, follows sound design principles, is easy to read, and has complete and meaningful documentation. Still, assessors can have different opinions on what makes a submission great, such as those that get awarded an A+ and exceptional submissions or those that are awarded an A++.
The rubric, which was provided to the students when they were completing their coursework, is a generalised rubric for each of the four assignments in the introductory Java course at our institution. It describes each of the four aspects, correctness, code elegance, documentation and readability, for the grades A+ - A++, A, B, C, D and F. While the +/- grades were not included in the rubric, we chose to include them in our study, as they were used when grading the original submissions.
@@
Some aspects of the rubric are very clear, especially for the grades at either end of the grade spectrum. For example, for correctness, an F is defined as “Significant details of a task are violated, or the program often exhibits incorrect behaviour.’ and an A+ - A++ is defined as “The application works as described in the assignment; all base tasks are completed; the implementation demonstrates originality, creativity and technical excellence…”.
@@
Using multiple graders is an inconsistent method of assessing programming assignments, especially for subjective aspects, such as code elegance, readability and documentation. None of our seven groups achieved even a tentative acceptable $\alpha>0.667$, with all being substantially lower with the highest $\alpha=0.45$ for correctness, $\alpha=0.19$ for code elegance, $\alpha=0.21$ for documentation, and $\alpha=0.17$ for readability.
There is likely to be less disagreement at either end of the grading spectrum, an F or an A+/A++, though rarely do any of these have complete agreement unless the submissions do not compile or do not contain any student code, with most groups being within one or two grades at these grades. Whereas the middle grades are not graded consistently, with most groups having a grade difference of three or more.
@@ 5.2. Are Individual Graders Consistent at Applying a Rubric? (RQ2)
During the post-study interview, 24 out of 28 graders stated that they graded consistently. However, our results show that when grading a repeated assignment at a later period, only two participants graded the duplicated assignment the same across all skills; in the post-study interview, participant 10 indicated that they noticed the duplicate. Both times, Participant 3 graded the duplicated assignment; they awarded all four skills an A++. They also gave identical feedback and said for correctness, “All listed core tasks completed… Challenge Tasks: Three challenge tasks are completed… demonstrating exceptional technical aptitude…”. While they gave identical feedback in this case, it does not indicate that they noticed the assignment was a duplicate, as they used a common grading practice of using the same feedback across multiple assignments. The average difference between grades across all skills was 1.4911, with correctness being the least consistently graded, with an average grade difference of 1.7857.
Correctness, being the most inconsistent when evaluating an individual’s consistency, is the inverse of evaluating consistency between multiple graders, where the subjectivness of code quality has a more prominent impact. The higher consistency of code elegance, readability, and documentation could be due to our participants being experienced programmers and having learnt programming standards in these areas, which they apply in their grading. Many of our participants have not graded before, and while some may have completed the assignment as a student, they would have only seen a limited set of other submissions. This could explain the variance in the grades, as many of our participants learnt how to judge the differing qualities of work, as they saw varying examples of high or low-quality submissions.
The environment in which the grading took place, the mood the graders were in, and the time of day in which the grading took place may have also impacted the consistency of the graders. Our participants felt like the environment they graded did not negatively impact their consistency, as most of them graded in the office or in silence at home. During our study, one participant was abroad and staying in a hotel; they said, “…we had visitors quite often, and so it was very hard to find a quiet space in my hotel room..”.
Some participants thought their mood could have affected their consistency, especially when they felt frustrated, bored or stressed. Participant 21 explained “…during the first set of assignments that we were grading when I had to submit a lot of submissions for my own semester as well. So I had a lot of hectic stress from that side.” Participant 28 noticed when they were getting frustrated or bored with the grading and took breaks to allow them to refocus. Participant 8 attempted to grade in a good mood and explained that “… if I am in a bad mood then I was complaining a lot…” and “…try to make empathy with the students as they are first-year students.”.
The time of day participants graded may have affected their consistency, as someone’s most productive period can vary from person to person. Participant 5 stated, “I think time of day affected me the most [out of the environment, mood and time of day] because I’m more of an afternoon person”. Whereas, Participant 8 said, “Evening is less desirable for me, as I am more of a morning person”.
Many factors can affect an individual consistency when awarding grades for assessment. Whether that be the graders’ experience with the assignment, their programming experience, the time of day they graded, the environment they graded in or the mood they were in when graded could have impacted their consistency.
The graders’ experience impacts consistency most, as they have learnt standards and formed their own opinions on what makes code elegant, readable and well-documented. As these standards and opinions are generic for any assignment, the participants were more consistent with grading these aspects than grading correctness, where they have to learn the specificity of what makes a submission good for this specific assignment.
While the environment the graders were in, the time of day they graded or the mood they were in could all impact the consistency of their grading, they could all be easily mitigated by grading in the same place, at the same time, and recognising when the graders are not in the right mood to grade. However, these mitigations are not always feasible, as grading deadlines, teaching and academic commitments can all factor into when, where and the mood they are in when they grade.
@@
Having more grading experience can lead to being more internally consistent when grading similar work. However, grading experience alone does not completely mitigate errors in consistency with experienced graders, especially when assessing subjective aspects, like documentation and readability. While summative assignments should be assessed by experienced graders, this is not always possible; a course may have an early-stage academic leading the course and assessment and preferred and experienced graduate teaching assistants may have graduated and left the institution and new graduate teaching assistants are asked to grade. These are just some factors where assignments may have to be assessed by inexperienced individuals.
@@ 7. Conclusion
We also captured where, when, and what mood the graders were in when they graded to investigate whether these factors affected the grader’s consistency. These factors were also discussed in a post-study interview, where we asked the participants if they noticed the duplicated submission. As our participants mostly graded in the office or at home, the environment had little impact on the consistency. However, our participants were more likely to report being impacted by the time of day they graded or the mood they were in while they were grading, as the time of day someone is productive varies from person to person. Many factors can affect a grader’s mood and grading, including other deadlines, frustration or boredom.
7.1. Implications for Human Grading
We have found that, in our study, typical graders are very inconsistent with each other and also are inconsistent with themselves. This suggests that current grading practices that use multiple assessors are likely to be quite unfair, which can lead to unfair outcomes for the students and demotivation among the students. Borela and Roy (2023) investigated using a workshop to train students in grading, and it may be that this is a necessary step to ensure fairness in grading.
We believe that our result warrants further consideration and investigation. Programming assignments are supposed to be relatively objectively judgeable and yet our research suggests that even with a shared rubric, inconsistency is the default outcome. This also lends support to the idea of investigating alternative grading approaches, such as those suggested by Decker et al. (2024) for making grading more equitable. They propose changing the grading scales (such as making them coarser) in order to speed up grading and reduce disputes about exact grading scores – although our study already uses a coarser scale than 0–100 and found inconsistency even on a coarser scale.
7.2. Implications for Automated Grading

### v2

@@ 5.3.1. Inter-Rater Metrics
Figure 6 shows the $\alpha$ for each of the seven groups and each skill. These results show that in our study, consistency was low, but were slightly more likely to agree on the grades for correctness and code elegance but did not agree on the grades for readability or documentation, with the average $\alpha$ being 0.2202, 0.0986, 0.0456, and 0.0548, respectively.
Figure 6 shows the results of Gwet’s $AC_{2}$. Following the same bounds defined by Krippendorf, no groups achieve acceptable agreement ($AC_{2}>0.800$) for any of the skills evaluated. However, some groups show partial agreement with $AC_{2}\geq 0.667$ for specific skills, though on average, no skill had an $AC_{2}\geq 0.667$, with the average $AC_{2}$ for correctness being 0.6286, for code elegance being 0.5924, for readability being 0.6383 and for documentation being 0.5966.
Out of the four skills that the participants graded, correctness was most likely to be graded more consistently, with five of the seven groups achieving a higher $\alpha$ than the other four skills, with groups 5-8 and 25-28 achieving the second most consistently graded skill for their respective groups. In comparison, all participant groups were similarly inconsistent when grading code elegance, readability and documentation, all skills achieving an average $\alpha<0.1$.
@@
5.3.2. Grading Experience
Group 17-20, which had the highest average $\alpha=0.2571$ consisted of three master’s students and one third-year student. Another group with a similar composition was group 1-4, consisting of three master’s students and one PhD student; however, they had an average $\alpha=0.1225$, the fourth lowest average $\alpha$.
5.3.3. Grade Range
To further explore the inter-rater reliability, we evaluated the spread of grades awarded by each group. Figure 8 shows the minimum, maximum, and mean correctness grades for group 17-20, which was the most consistent when grading correctness, with an $\alpha=0.45$, and for group 5-8, which was the least consistent when grading correctness, with an $\alpha=-0.069$. Appendix C shows the distribution of grades for all groups and all grades.
For our most consistent group (group 17-20), only assignments awarded an F (assignments 80 and 98) – since they either did not compile or did not include any student code – were assessed with complete agreement. The original authors of assignment 80 submitted the template code they provided, and assignment 98 did not compile or, if the assessors fixed the compilation error, had an unhandled NumberFormatException, causing the simulator not to execute.
@@ 5.3.3. Grade Range
Figure 8. These figures are for the highest and lowest $\alpha$ and show the minimum, maximum and mean correctness grade awarded by the participants for each assignment. Plots for all groups and all grades can be found in Appendix C.
To further investigate the reliability when awarding each grade, we stipulate that assignments grade ranges less than or equal to two (e.g. A to B+ is two grades away, and A to A- is one grade away) have good reliability; those with a range of three (e.g. A+ to B+) have mediocre reliability, and those with a range greater than three (e.g. A- to C+) have poor reliability; where the grade range is the maximum awarded grade minus the minimum awarded grade for each assignment.
Figure 9 shows the proportion of the reliability for each grade and skill and shows that our participants were more reliable at awarding F and D- grades across all four skills than any other, with the middle grades (C- to B+) being the most unreliably graded across all four skills. The higher grades (A- to A++) were reliable or partiality reliable in less than 20% of assignments graded across all four skills.
@@ 5.3.4. An In-Depth Analysis of the Grades Awarded by Group 17-20
This trend of a few assignments being somewhat consistently graded and many being inconsistently graded holds for readability and documentation, with Appendix C showing the grade range for each skill. Figure 9 shows this trend for each grade, with the middle grades (D - B+) more likely to have a larger variance in awarded grades across all four skills. While the middle grades showed no cases of any grading achieving acceptable levels of consistency, the high (A++ - A) and low ranges (F - D-) displayed at least some instances of grader agreement. Grades F and D- have the largest portions of grade ranges, being less than or equal to two. At the other end of the grade spectrum, grades A, A+ and A++ have a smaller proportion of grade ranges that are less than or equal to two.
5.4. Grading Consistency of Individual Graders (RQ2)
5.4.1. Grade Range
5.4.2. Post-Study Interview
During the post-study interview, 24 out of 28 graders stated that they graded consistently. However, our results show that when grading a repeated assignment at a later period, only two participants graded the duplicated assignment the same across all skills. Only two participants, three and ten, provided the same grade in both batches for all four skills. Participant 10 noticed the duplicate, and during their post-study interview, participant 10 mentioned that they noticed the duplicate and changed the grades for both batches to F as they believed it was plagiarism, while the other five participants followed the instruction that if you notice a duplicate, treat it as you are grading it the first time; as this was a pair programming assignment, duplicates could still exist within the dataset. Participant 3 did not notice the duplicate and said during their post-study interview, “…like some assignments used the same animals… but [the implementation] was a little bit different so not like an exact copy of those two assignments…”.
5.4.3. Grading Experience
Figure 11 shows the distribution by grading experience of the absolute distance between grades awarded in batch one and batch two for each of the four skills. Those with no prior grading experience have the largest difference between grades across all skills. Those with 1-2 years of grading experience have at most a difference of three grades between the first and second time they graded the duplicated assignment, and those with one or more years of grading have a grade difference of less than one for most skills. However, readability and documentation for those with two or three years of grading experience had a maximum grade distance of two or three, respectively.
6. Discussion
6.1. How Consistently Does a Group of Graders Apply a Rubric? (RQ1)
Consistency was very poor overall. Even the best values achieved for Krippendorf’s $\alpha$ reflect distinctly poor consistency, with none of our seven groups of graders achieving an $\alpha>0.667$, the lower bound to consider reliable (Krippendorff, 2013), with the average $\alpha=0.10480$ across all groups and skills.
While the correctness of a program is, in most cases, inherently objective, there can be elements of subjectiveness, especially within open-ended coursework assignments. The nature of open-ended assignments, where students decide how they implement a set of features using their ideas and creativity, can make it difficult for the groups of assessors to grade consistently. However, open-ended assignments allow instructors to design assignments that allow students to explore their creativity, which is a key educational goal (Chan and Yuen, 2014; Richardson and Mishra, 2018; Wagner, 2010), mimic professional development experiences, and improve students’ self-efficacy in programming (Sharmin et al., 2019).
Furthermore, the subjectiveness of open-ended programming assignments extends not just to whether the code is correct and meets the requirements but to the quality of the code itself (Kirk et al., 2024a). While some aspects of code quality are shared, such as consistent indentation and meaningful variable names, the nuances of programming style and design are individualised and remain active research topics not just within computer science education (Stegeman et al., 2016; Kirk et al., 2024b), but within professional practice as well (Liang et al., 2023).
While having experienced graders may increase the consistency of assessment, this is not always possible as the typical demographics of teaching assistants are undergraduate, master’s or PhD students (Mirza et al., 2019; Riese and Bälter, 2022; Riese et al., 2021; Wald and Harland, 2018), and many are new to assessment and have little experience; which is the case in our study. Having those with many years of grading experience is typically rare, as undergraduate teaching assistants are typically late-stage and are about to graduate, and graduate teaching assistants within Europe typically graduate within three to four years.
However, those with one year of grading experience had similar inter-rater reliability to those with no previous grading experience, indicating that more grading experience alone will not improve the grading consistency to an acceptable level. This supports the idea that to conduct a consistent assessment, assessors require adequate training to interpret and apply a rubric (Riese and Bälter, 2022; Kristiansen et al., 2024), and that assessment is complex and intuitive (Bloxham et al., 2016; Anders Jönsson and Hartell, 2021).
Rubric design plays an important role in the validity and consistency of the rubric (Jonsson and Svingby, 2007). The rubric provided to our participants follows the standard practice for rubric design (Andrade, 2005; Reddy and Andrade, 2010) and includes evaluation criteria, the definitions of quality and a scoring strategy.
@@ 6.1. How Consistently Does a Group of Graders Apply a Rubric? (RQ1)
However, as with the subjectiveness of programming, especially with open-ended assessment and code quality, as we have previously discussed, aspects of the rubric are subjective. Some aspects of the rubric are very clear, especially for the grades at either end of the grade spectrum. For example, for correctness, an F is defined as “Significant details of a task are violated, or the program often exhibits incorrect behaviour.’ and an A+ - A++ is defined as “The application works as described in the assignment; all base tasks are completed; the implementation demonstrates originality, creativity and technical excellence…”.
@@
There is likely to be less disagreement at either end of the grading spectrum, an F or an A+/A++, though rarely do any of these have complete agreement unless the submissions do not compile or do not contain any student code, with most groups being within one or two grades at these grades. Meanwhile, the middle grades are not graded consistently, with most groups having a grade difference of three or more.
While a more explicit rubric may improve the consistency of the grading, it would have reduced the validity of the assignment (Jonsson and Svingby, 2007), as a more explicit rubric would require a less open and creative assignment, limiting the students’ self-efficacy (Sharmin et al., 2019), and learning opportunities. Furthermore, even an explicit and analytic rubric may not infer consistent assessment, as grading consistently can be extremely difficult even with well-designed assessment criteria (Anders Jönsson and Hartell, 2021).
@@ 6.2. Are Individual Graders Consistent at Applying a Rubric? (RQ2)
Compared to the consistency of multiple graders, readability, code elegance, and documentation achieved a higher individual consistency. The higher consistency of code elegance, readability, and documentation could be due to our participants being experienced programmers and having a well-defined internal preference for code elegance, readability and documentation (Mohan and Gold, 2004; Li et al., 2018).
The lower consistency of correctness could be due to many of our participants having not graded before, which is typical when using teaching assistants (Mirza et al., 2019; Riese and Bälter, 2022; Riese et al., 2021; Wald and Harland, 2018). While some may have completed the assignment as a student, they would have only seen a limited set of other submissions. Another aspect could be the participants having to learn the specificity of what makes a submission good for this specific assignment; as the assignment is open-ended many different solutions can meet the requirements to the same standard, but with the specificity of the rubric, that standard is up interpretation by the grader. Similar to multiple grader consistency, a more specific rubric could increase the consistency, but at the cost of the validity (Jonsson and Svingby, 2007) and creativity of the assignment.
@@
Having more grading experience can lead to being more internally consistent when grading similar work, and supports the emphasis on providing appropriate training for teaching assistants (Kristiansen et al., 2024; Wald and Harland, 2018). However, grading experience alone does not completely mitigate errors in consistency with experienced graders (Riese and Bälter, 2022), especially when assessing subjective aspects like documentation and readability.
6.3. Implications for Human Grading
We have found that, in our study, typical graders are very inconsistent with each other and also are inconsistent with themselves. This suggests that current grading practices that use multiple assessors are likely to be quite unfair, which can lead to unfair outcomes for the students and demotivation among the students. Many suggest the importance of training assessors (Kristiansen et al., 2024; Wald and Harland, 2018; Borela and Roy, 2023), and it may be that this is a necessary step to ensure fairness in grading.
We believe that our result warrants further consideration and investigation. Programming assignments are supposed to be relatively objectively judgeable and yet our research suggests that even with a shared rubric, inconsistency is the default outcome, which supports the idea that grading is ‘complex, intuitive and tacit’ (Bloxham et al., 2016). This also lends support to the idea of investigating alternative grading approaches, such as those suggested by Decker et al. (2024) for making grading more equitable. They propose changing the grading scales (such as making them coarser) in order to speed up grading and reduce disputes about exact grading scores – although our study already uses a coarser scale than 0–100 and found inconsistency even on a coarser scale.
6.4. Implications for Automated Grading
@@ 8. Conclusion
Our study supports the notion that assessment is complex and intuitive (Bloxham et al., 2016), and that having a well-designed open-ended assignment may not imply consistent evaluation (Jonsson and Svingby, 2007), and suggests that there is a balance to be found between the openness of the assignment and the specificity of the rubric. It also highlights that further research is required to improve the consistency of a group of graders and individual graders when grading open-ended programming assignments, whether it be alternative grading practices or methods of effectively training undergraduate and graduate teaching assistants.